OpenStudy (anonymous):

Which ordered pairs from the replacement set are solutions to the equation? y = -2x + 9 ; Replacement Set {((-1, 11), (0, 8), (1, 7), (2, 5)} Answer (-1, 11) (0, 8) (1, 7) (2, 5)

OpenStudy (anonymous):

First, recognize that in each point, the first number is the x-coordinate and the second number is the y-coordinate. For example, if we choose (2, 5), then we have x=2 when y=5. Substitute these into the equation: y = -2x + 9 5 = -2(2) + 9 5 = -4 + 9 5 = 5 Since we got a true statement, it means that (2, 5) satisfies the equation (if we had gotten something weird like 5 = 11, then that point would NOT satisfy the equation). You need to go through each pair of coordinates to figure out which ones work, because there are several that work.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Just finished checking them all! You made that so easy! Thanks!

OpenStudy (anonymous):

No problem :) .